# Evolution of a Long-Lived Deep-Seated Main-Sequence Magnetic Field During White Dwarf Cooling

### v1

@@ 4 Comparison with observations
The inclusion of models with d and d+q geometry sets an error in the estimation of $B_{\mathrm{MS}}$. Depending on the cooling age of the WD, having a d+q geometry implies that the $B_{\mathrm{MS}}$ prediction is larger than the d case at most by a factor of $\sim 2$. As mentioned in Section 3.1, the d+q geometry can better reproduce the off-centered dipole nature that fits many MWDs (Hardy et al., 2023; Moss et al., 2025), although the individual contributions of dipole and quadrupole components to the d+q model may depend on each MWD, and the initial conditions assumed here only set an approximate estimation for how this model is different from the simple dipole. While including a turbulent diffusion component only sets an error estimation for those MWDs that are already undergoing crystallization, for the oldest ones, the $B_{\mathrm{MS}}$ prediction can vary up to a factor of $\sim 4$ when contrasting the d+q geometry including turbulent diffusion with the d geometry without turbulent transport.
@@
For a comparison of the magnetic field strength observed in MWDs and the predicted from our models, we considered the magnetic field at the start of the WD phase $B_{0}$, to be the one predicted from magnetic flux conservation of the equipartition prediction of $B_{\mathrm{MS}}$, and considered the diffusion for about 9 Gyr. In Figure 7, we show the evolution of $B_{0}$ for the models with masses $0.6\ M_{\odot}$ to $1\ M_{\odot}$, considering the d and d+q geometries and turbulent diffusion and not. We compare these models with the observed fields of MWDs in a 100 pc sample of the MWDD with masses in the range $0.55$-$1.05\ M_{\odot}$ in bins of mass separated by $0.1\ M_{\odot}$, as shown for different panels of Figure 7. We note that the dipole model without turbulent diffusion appears almost as an upper limit for the observations for each mass. Additionally, we note that many of the observations seem to roughly follow the pattern given by the models: the magnetic field strength is increasing with time almost saturating after the model reaches a maximum, again confirming that the late appearance of MWDs can still be explained by a main-sequence dynamo (Camisassa et al., 2024).
@@
From Figures 4 and 5 it can be inferred that, if $B_{\mathrm{0}}$ was almost the same for any WD mass, the observed magnetic fields would be weaker on average for less massive WDs due to what we predict from solving the evolution of the magnetic field. This behavior is even more prominent in the models in Figure 7 because the equipartition prediction increases with the WD mass, and because the factor of Radius reduction at the magnetic boundary, $\frac{R_{\mathrm{MS,b}}}{R_{\mathrm{WD,b}}}$, is larger for more massive stars. In the lower right panel of Figure 7 we compare the range of the observed magnetic field strengths per mass bin; for this, the error bars in the vertical axis cover the whole field strength for the given bin and have as a data point the logarithmic average of the magnetic fields in the bin. On top of each bin we plot the model range, where the upper limit is given by the maximum field strength predicted by equipartition and the arrow is at the minimum model value predicted for the age of the youngest WD in the bin. We see that the average magnetic field strengths in this 100 pc sample increase with mass, which is similar to what we predict with the evolution models.
@@
It has been reported in the literature that MWDs are on average more massive than non-MWDs (Ferrario et al., 2020). Whether more massive MWDs have stronger magnetic fields is still poorly constrained. Recently, Moss et al. (2025) identified 167 MWDs, with 87 new discoveries, within the 100 pc sample of Kilic et al. (2025). In this new MWDs sample, it is reported that at least two groups are identified in a three-parameter space analysis, one group of older ($\sim 2.9$ Gyr), less massive ($\sim 0.7\ M_{\odot}$), and with weaker magnetic fields ($\sim 3.7$ MG); and other group of younger ($\sim 1.8$ Gyr), more massive ($\sim 0.96\ M_{\odot}$), and with stronger fields ($\sim 84$ MG). These results favor our analysis, since our models predict that stronger magnetic fields tend to appear earlier in more massive WDs.
@@
Additionally, Moss et al. (2025) noted that more MWDs in their sample can be explained by the emergence of a main-sequence magnetic field emerging during the WD phase than by the magnetic field generated through a crystallization-driven dynamo. However, they only used the breakout times estimated by Camisassa et al. (2024) and Blatman & Ginzburg (2024a), and the strength of the magnetic field was not addressed to compare the theories. Here we show how adding the magnetic field strength in the calculation may also favor the appearance of a main-sequence magnetic field during the WD phase to explain some magnetic MWDs of any mass. Conversely, Castro-Tapia et al. (2024b) predicted that the crystallization-driven dynamo theory can only explain MWDs of up to a few MG in the best-case scenario.
5 Summary and Conclusions
@@
We analyze two different geometries for the poloidal field: a simple dipole and a dipole plus a quadrupole component. We find that the latter is consistent with an off-centered dipole model, which has been proposed to explain photometric and spectroscopic data of MWDs (Hardy et al., 2023). In doing so, we observe that the d+q geometry yields weaker predictions for the surface magnetic field strengths for a constant initial magnetic field $B_{0}$ in the WD phase. This occurs due to the more rapid decay of the quadrupole component in the magnetic field expansion (e.g. Cumming, 2002). Additionally, we also consider models with turbulent diffusion during crystallization as in Castro-Tapia et al. (2024b). Having these different models lead to surface magnetic field strengths that vary by a factor of 2 to 4 for the same cooling age and mass.
@@
By taking a sample limited to 100 pc in the Montreal WD Database in the range of 0.55 to 1.05 $M_{\odot}$ and the prediction of $B_{0}$ from equipartition for our evolution models, we compare the field strength as a function of time for fixed masses. We find that the models with different geometries and with and without turbulent diffusion cover a range that can fit many of the strengths observed in MWDs. However, in general, using the equipartition assumption for $B_{0}$ appears almost as an upper limit for the field prediction at different masses, which would be expected if the magnetic field given by the core-convective dynamo is partially diffused during stellar evolution. Mohapatra & Blackman (2024) recently pointed out that the core-convective dynamo may be unlikely if turbulent diffusion acts to rapidly diffuse the magnetic field when a stage of no magnetic field generation occurs during convection. They estimate the time for turbulent diffusion to be only $\gtrsim 10$ yr, but only assumed that the field would dissipate in the convection zone. Conversely, Hidalgo et al. (2024) showed that only a small fraction of the magnetic energy can reach the surface as diffusion acts in the region outside the convective core (even with similar turbulent diffusion times); then, more detailed MHD simulations of convection in the different stellar evolution stages are necessary to understand the survival and evolution of the core-convective generated magnetic field.
@@
We also follow up on the prediction of Camisassa et al. (2024) and show that, with a more detailed evolution of the magnetic field, the late appearance of high magnetic field strengths is not necessarily related to crystallization and may be explained almost solely by the appearance of a deep-seated magnetic field that was generated during the main-sequence and survive until the WD phase. Furthermore, such theory may also explain MWDs with masses $\lesssim 1.1\ M_{\odot}$ that appear before the crystallization onset, which, in the limited 100 pc that we use, tend to be weaker than those that appear after crystallization. However, note that this behavior is better explained if we take the cooling age at which our models reach a maximum surface magnetic strength value for each mass (Figure 7). The observed magnetic fields that are younger than the age at which the maximum model magnetic field is reached tend to be weaker than those that are older than that point.

### v2

@@ IV Comparison with observations
The inclusion of models with d and d+q geometry sets an error in the estimation of $B_{\mathrm{MS}}$. Depending on the cooling age of the WD, having a d+q geometry implies that the $B_{\mathrm{MS}}$ prediction is larger than the d case at most by a factor of $\sim 2$. As mentioned in Section III.1, the d+q geometry can reproduce the off-centered dipole nature that fits many MWDs (Hardy et al., 2023; Moss et al., 2025), although the individual contributions of dipole and quadrupole components to the d+q model may depend on each MWD, and the initial conditions assumed here only set an approximate estimation for how this model is different from the simple dipole. While including a turbulent diffusion component only sets an error estimation for those MWDs that are already undergoing crystallization, for the oldest ones, the $B_{\mathrm{MS}}$ prediction can vary up to a factor of $\sim 4$ when contrasting the d+q geometry including turbulent diffusion with the d geometry without turbulent transport.
@@
For a comparison of the magnetic field strength observed in MWDs and the predicted from our models, we considered the magnetic field at the start of the WD phase $B_{0}$, to be the one predicted from magnetic flux conservation of the equipartition prediction of $B_{\mathrm{MS}}$, and considered the diffusion for about 9 Gyr. In Figure 7, we show the evolution of $B_{0}$ for the models with masses $0.6\ M_{\odot}$ to $1\ M_{\odot}$, considering the d and d+q geometries and turbulent diffusion and not. We compare these models with the observed fields of MWDs in a 100 pc sample with masses in the range $0.55$-$1.05\ M_{\odot}$ in bins of mass separated by $0.1\ M_{\odot}$, as shown for different panels of Figure 7. We note that the dipole model without turbulent diffusion appears almost as an upper limit for the observations for each mass. Additionally, we note that many of the observations seem to roughly follow the pattern given by the models: the magnetic field strength is increasing with time, almost saturating after the model reaches a maximum, again confirming that the late appearance of MWDs can still be explained by a main-sequence dynamo (Camisassa et al., 2024).
In principle, factors such as the rotation of the star should affect the scaling of the magnetic field when it is generated during the main-sequence. However, numerical simulations on convective dynamos show that the magnetic field can reach a significant fraction of equipartition (Hidalgo et al., 2024), then it seems reasonable considering such a scaling as an upper limit. Other uncertainty sources in determining the magnetic boundary are related to the whole stellar evolution history (Camisassa et al., 2024), such as convective boundary determinations, mass loss rates during the giant phases, and the differences in the initial-to-final-mass relation when assuming different stellar evolution models (Cummings et al., 2018). While we did not attempt to explore these uncertainties, it is worth mentioning that our theory is presented in an idealistic scenario, where the magnetic field survives almost intact inside the radiative zones of the star during the giant phases.
@@
From Figure 4 it can be inferred that, if $B_{\mathrm{0}}$ was almost the same for any WD mass, the observed magnetic fields would be weaker on average for less massive WDs due to what we predict from solving the evolution of the magnetic field. This behavior is even more prominent in the models in Figure 5 and 7 because the equipartition prediction increases with the WD mass, and because the factor of radius reduction at the magnetic boundary, ${R_{\mathrm{MS,b}}}/{R_{\mathrm{WD,b}}}$, is larger for more massive stars. In the lower right panel of Figure 7, we compare the observed magnetic field strengths at each mass; for this, we show the data in the whole mass range of the combined samples of MWDs (black dots) and a data point for the logarithmic average of the magnetic fields in the bins used in the previous panels. On top of each bin, we plot the model range, where the upper limit is given by the maximum field strength predicted, and the arrow is at the minimum model value predicted for the age of the youngest WD in the bin. We see that the average magnetic field strengths in this 100 pc sample increase with mass, which is similar to what we predict with the evolution models.
As a more appropriate test of the apparent correlation between magnetic field strength and WD mass, we made a linear fit in the logarithmic space of mass and $B$ for all the data points shown (in black) of the combined samples, which means that we fit an exponential relation characterized by $B\propto M^{a\pm\sigma_{a}}_{\mathrm{WD}}$. We find a value $a=4.592\pm 0.678$, showing a positive correlation between the mass and $B$, which is true for our fit even if we consider the case where the exponential factor is $a-5\sigma_{a}$. In the lower right panel of Figure 7, we also show the fit made (magenta line) considering the $\pm 5\sigma$ range of the fitted parameters (shaded magenta area). This result shows that despite the large dispersion in the data, the positive correlation between WD mass and magnetic field strength might be significant.
@@
It has been reported in the literature that MWDs are, on average, more massive than non-MWDs (Ferrario et al., 2020). Whether more massive MWDs have stronger magnetic fields is still poorly constrained. Recently, Moss et al. (2025) identified 167 MWDs, with 87 new discoveries, within the 100 pc sample of Kilic et al. (2025). In this new MWDs sample, it is reported that at least two groups are identified in a three-parameter space analysis, one group of older ($\sim 2.9$ Gyr), less massive ($\sim 0.7\ M_{\odot}$), and with weaker magnetic fields ($\sim 3.7$ MG); and other group of younger ($\sim 1.8$ Gyr), more massive ($\sim 0.96\ M_{\odot}$), and with stronger fields ($\sim 84$ MG). These results favor our analysis, since our models predict that stronger magnetic fields tend to appear earlier in more massive WDs.
@@
Additionally, Moss et al. (2025) noted that more MWDs in their sample can be explained by the emergence of a main-sequence magnetic field during the WD phase than by the magnetic field generated through a crystallization-driven dynamo. However, they only used the breakout times estimated by Camisassa et al. (2024) and Blatman & Ginzburg (2024a), and the strength of the magnetic field was not addressed to compare the theories. Here, we show how adding the magnetic field strength in the calculation may also favor the appearance of a main-sequence magnetic field during the WD phase to explain some magnetic MWDs of any mass. Conversely, Castro-Tapia et al. (2024b) predicted that the crystallization-driven dynamo theory can only explain MWDs of up to a few MG in the best-case scenario.
V Summary and Conclusions
@@
We analyze two different geometries for the poloidal field: a simple dipole and a dipole plus a quadrupole component. We find that the latter is naturally consistent with an off-centered dipole model, which has been proposed to explain photometric and spectroscopic data of MWDs (Hardy et al., 2023). In doing so, we observe that the d+q geometry yields weaker predictions for the surface magnetic field strengths for a constant initial magnetic field $B_{0}$ in the WD phase. This occurs due to the more rapid decay of the quadrupole component in the magnetic field expansion (e.g. Cumming, 2002). Additionally, we also consider models with turbulent diffusion during crystallization as in Castro-Tapia et al. (2024b). Having these different models lead to surface magnetic field strengths that vary by a factor of 2 to 4 for the same cooling age and mass.
@@
By taking a sample limited to 100 pc in the Montreal WD Database, combined with the MWDs in Bagnulo & Landstreet (2022), in the range of 0.55 to 1.05 $M_{\odot}$ and the prediction of $B_{0}$ from equipartition for our evolution models, we compare the field strength as a function of time for fixed masses. We find that the models with different geometries and with and without turbulent diffusion cover a range that can fit many of the strengths observed in MWDs. However, in general, using the equipartition assumption for $B_{0}$ appears almost as an upper limit for the field prediction at different masses, which would be expected if the magnetic field given by the core-convective dynamo is partially diffused during stellar evolution. Mohapatra & Blackman (2024) recently pointed out that the core-convective dynamo may be unlikely if turbulent diffusion acts to rapidly diffuse the magnetic field when a stage of no magnetic field generation occurs during convection. They estimate the time for turbulent diffusion to be only $\gtrsim 10$ yr, but only assumed that the field would dissipate in the convection zone. Conversely, Hidalgo et al. (2024) showed that only a small fraction of the magnetic energy can reach the surface as diffusion acts in the region outside the convective core (even with similar turbulent diffusion times); then, more detailed MHD simulations of convection in the different stellar evolution stages are necessary to understand the survival and evolution of the core-convective generated magnetic field.
@@
We also follow up on the prediction of Camisassa et al. (2024) and show that, with a more detailed evolution of the magnetic field, the late appearance of high magnetic field strengths is not necessarily related to crystallization and may be explained almost solely by the appearance of a deep-seated magnetic field that was generated during the main-sequence and survived until the WD phase. Furthermore, such a theory may also explain MWDs with masses $\lesssim 1.1\ M_{\odot}$ that appear before the crystallization onset, which, in the limited 100 pc that we use, tend to be weaker than those that appear after crystallization. However, note that this behavior is better explained if we take the cooling age at which our models reach a maximum surface magnetic strength value for each mass (Figure 7). The observed magnetic fields that are younger than the age at which the maximum model magnetic field is reached tend to be weaker than those that are older than that point.